OpenStudy (anonymous):

I dont understand this at all,im so sorry someone help! What is the value of the expression –12 – 3 + 9?

OpenStudy (anonymous):

its -6

OpenStudy (anonymous):

-12 - 3 = -15 -15 + 9 = -6

OpenStudy (lncognlto):

Yep, that's correct. ^-^